DiabolikLover:

if a newton is kg/(m/s^2) simplify the units

ramen:

Are you sure this is the full question?

DiabolikLover:

yes, i know my teacher makes no sense. it has to deal with force and weight

jimthompson5910:

Hint: \(\large \frac{A/B}{C/D} = \frac{A}{B}*\frac{D}{C} = \frac{AD}{BC}\) Also, it might help to think of "kg" as "kg/1"